# SR-Scientist: Scientific Equation Discovery With Agentic AI

### v1

@@ Abstract
[Uncaptioned image] Code: https://github.com/GAIR-NLP/SR-Scientist[Uncaptioned image] Models: SR-Scientist-30B
Recently, Large Language Models (LLMs) have been applied to scientific equation discovery, leveraging their embedded scientific knowledge for hypothesis generation. However, current methods typically confine LLMs to the role of an equation proposer within search algorithms like genetic programming. In this paper, we present SR-Scientist, a framework that elevates the LLM from a simple equation proposer to an autonomous AI scientist that writes code to analyze data, implements the equation as code, submits it for evaluation, and optimizes the equation based on experimental feedback. Specifically, we wrap the code interpreter into a set of tools for data analysis and equation evaluation. The agent is instructed to optimize the equation by utilizing these tools over a long horizon with minimal human-defined pipelines. Empirical results show that SR-Scientist outperforms baseline methods by an absolute margin of 6% to 35% on datasets covering four science disciplines. Additionally, we demonstrate our method’s robustness to noise, the generalization of the discovered equations to out-of-domain data, and their symbolic accuracy. Furthermore, we develop an end-to-end reinforcement learning framework to enhance the agent’s capabilities.
@@ 1 Introduction
In the era of Agentic AI, Large Language Models (LLMs) have evolved from simple knowledge retrievers to agentic models capable of completing complex tasks by interacting with their environments (Schneider, 2025), as exemplified by products like Claude Code (Anthropic, 2025) and Gemini CLI (Google, 2025). These agents exhibit many characteristics of human scientists, such as engaging in long-horizon interaction with environmental feedback, operating with autonomy, and relying less on predefined pipelines (Newell et al., 1972). However, in most current work using LLMs for scientific discovery (Shojaee et al., 2024; Novikov et al., 2025; Romera-Paredes et al., 2024), LLMs serve as static components within human-crafted pipelines, lacking the autonomy to generate and refine hypotheses through active environmental interaction. Building a scientific discovery framework around agentic models could therefore shift the paradigm from using LLMs as passive tools to empowering them as autonomous agents that drive the entire discovery lifecycle. In this paper, we focus on equation discovery, a fundamental task in science.
Mathematical equations are central to scientific progress, serving as concise and interpretable models of physical phenomena (Lemos et al., 2022; Batra et al., 2021; Hernandez et al., 2019). The task of data-driven equation discovery, also known as symbolic regression (SR), is an NP-hard problem due to its vast search space (Virgolin and Pissis, 2022). Traditional methods have relied on Genetic Programming (GP) for combinatorial search (Cranmer, 2023) or on deep neural networks trained on large-scale synthetic data for direct prediction (Kamienny et al., 2022; Biggio et al., 2021), which often suffer from efficiency and scalability issues. Recently, the research community has begun to embed LLMs into GP algorithms as equation proposers, leveraging their scientific prior knowledge to generate more effective hypotheses (Shojaee et al., 2024; Grayeli et al., 2024). While this can enhance search efficiency, LLMs primarily serve as equation generators in a fixed pipeline, which lacks the autonomy to gain insight for equation design through tools.
@@
To validate the effectiveness of our framework, we evaluate SR-Scientist using 5 different models as backbone LLMs on datasets covering 4 science disciplines carefully designed to prevent the LLMs from relying on memorized equations from their training data (Shojaee et al., 2025). Results demonstrate that SR-Scientist consistently outperforms state-of-the-art SR methods, with four of the models achieving an absolute margin of 6% to 35% over baselines. We also observe significant improvement after applying RL training. Additionally, we demonstrate our method’s robustness to noise, the generalization of the discovered equations to out-of-domain data, and their symbolic accuracy. Furthermore, our ablation studies highlight the importance of data analysis, the experience buffer, and long-horizon optimization.
@@ 2 Related Work
SR aims to identify an interpretable mathematical expression that accurately describes the relationship between variables in observed data (Makke and Chawla, 2024; Cava et al., 2021). Current popular methods can be categorized as follows: 1) GP-based Methods: This line of methods uses constrained representations, such as expression trees, to define a search space composed of symbolic operators, constants, and variables (Cranmer, 2023). Evolutionary or genetic programming is then applied to search for solutions. 2) Deep learning Methods: These methods leverage deep neural networks to learn a direct mapping from numerical observations to their underlying mathematical expressions through online (Petersen et al., 2021; Landajuela et al., 2022) or offline learning (Kamienny et al., 2022; Biggio et al., 2021). 3) LLM-based methods: Due to the efficiency and scalability issues of the above methods, the research community has integrated LLMs with traditional techniques like GP, leveraging the models’ extensive knowledge to generate more informed hypotheses (Shojaee et al., 2024; Grayeli et al., 2024; Ma et al., 2024; Wang et al., 2025b). While promising, current hybrid methods still lack the direct analysis of observed data through tools to gain insights. Furthermore, the LLM’s behavior is often fixed, lacking the autonomy to decide its actions in a goal-oriented manner. Moreover, most work focuses on using LLMs for inference, without exploring how these models evolve their abilities through methods like RL.
@@ 3.2 Inference Framework
Tool Design
We provide a code interpreter as the primary tool, enabling the agent to write code for data analysis and performance evaluation. Specifically, we wrap the code interpreter into two common tools for the tasks, data analyzer and equation evaluator, denoted as $T_{1}$ and $T_{2}$, respectively. For $T_{1}$, we link it to the observed data and include a code example in the prompt demonstrating how to access the data. Through this tool, the agent can write code to analyze the data, including printing the data, conducting statistical analysis, or analyzing the residuals between the predicted value and the true value. For $T_{2}$, following the approach of Shojaee et al. (2024), we design the tool to accept an equation skeleton with placeholders for constants in code format (See Figure 1 or 5 for the equation example). During the execution phase, the BFGS algorithm is implemented to optimize these constants and then report the performance of the equation. It also accepts equations with constants decided through other methods like data analysis. This tool prevents the agent from writing repetitive code for equation evaluation during the long-horizon exploration process.
@@ 3.3 Training Framework
Following the approach of Shojaee et al. (2025), we employ a mixed strategy of rule-based and model-based data synthesis. For each scientific scenario covering multiple variables, we instruct an LLM to synthesize potential relationships between the variables, typically in the form of an equation skeleton. For each skeleton, we use the model to determine the values for the constants based on their physical significance and thereby construct the full equation. Once the complete equation is formulated, we define an appropriate range of values for its variables and synthesize the observed data accordingly. This dataset is then split into training and evaluation sets. The training data is accessible to the agent during rollouts, while the evaluation data is used to measure performance and calculate the reward.
@@
where $s_{\text{max}}$ represents the maximum MAPE for which a non-zero reward can be gained. Additionally, we explore other reward functions and present the discussion in Appendix A.3.2.
@@ 4.1 Setup
Due to the vast corpus that LLMs have been pretrained on, the evaluation dataset must be carefully curated to prevent LLMs from having memorized the equations. We evaluate our methods on the synthetic part of the LLM-SRBench (Shojaee et al., 2025), denoted as LSR-Synth [^2] We also considered LSR-Transform (Shojaee et al., 2025) as a candidate for the evaluation dataset. However, we found that it might still have memorization issues, as nearly 50% of the problems achieved a sufficiently small error after the first iteration.. It combines known terms in the underlying equation with synthetic, novel terms to create problems that go beyond memorization. It contains 129 problems, spanning four scientific domains: chemistry (36), biology (24), physics (44), and material science (25). For each problem, it has three types of datasets: a training set accessible to the SR method, an in-domain (ID) test set, and an out-of-domain (OOD) test set.
@@
We use accuracy-to-tolerance as our main metric. We find it serves as a more robust metric when aggregating the overall results covering multiple problems than others, like Normalized Mean Squared Error. It is also more challenging than other metrics like $R^{2}$. Please refer to Appendix A.4 for a detailed analysis. Given a desired tolerance threshold $\tau$, the metric calculates whether the predicted values and the ground truth values satisfy it as follows:
@@
We discard the 5% worst predictions to prevent outliers due to the presence of the max operator, following Biggio et al. (2021); Kamienny et al. (2022). Additionally, we present symbolic accuracy, which checks if the discovered equations are identical to the ground truth equations. We present the details for the symbolic accuracy calculation in Appendix A.7.
@@
We comprehensively compare our methods against several baseline methods. For methods without LLMs, we include GPLearn, E2E (Kamienny et al., 2022), NeSymReS (Biggio et al., 2021), DSR (Petersen et al., 2021), uDSR (Landajuela et al., 2022), and PySR (Cranmer, 2023). For the LLM-based methods, we include LLM-SR and LaSR for comparison. For methods without LLMs, we constrain the number of equation candidates for each problem to 100,000. For methods with LLMs, we constrain the LLM calls to 1,000, with at most 1,000 equation candidates from each problem. For SR-Scientist, we set the maximum number of turns per iteration to 25, with a total of 40 iterations. We present the configuration details of the baseline methods in Appendix A.1 and the configuration of SR-Scientist in Appendix A.2. We evaluate our method with different models as backbone LLMs, including Qwen3-Coder-480B-A35B (Qwen, 2025), GLM-4.5-Air (Zeng et al., 2025a), GPT-OSS-120B (OpenAI, 2025), GPT-OSS-20B, and Qwen3-Coder-30B-A3B. Due to computation costs and training infrastructure limitations, we use only Qwen3-Coder-30B-A3B for RL training, while the other models are used for inference only. However, we also verify the effectiveness of RL training on other models and present the results in Appendix A.3.4. For each experiment, we repeat it three times and report the average value to reduce noise.
@@
For training data construction, we synthesize training data for the scientific scenarios in LSR-Synth. To prevent data contamination, two authors of this paper independently and carefully checked the equation skeletons to prevent duplicates and confirmed their agreement. Finally, we obtained 1024 problems for RL training. Please refer to Appendix A.3.1 for details. During rollout, we set the maximum number of turns to 20 and train for 60 steps. We set $s_{\text{max}}$ to 100% and $s_{\text{goal}}$ to 0.1% in the reward design. The detailed training configure can be found in the Appendix A.3.3.
@@ 4.2 Main Results
Table 1 presents the accuracy results. In terms of overall accuracy, SR-Scientist consistently outperforms the baseline methods, with four of the models achieving an absolute performance margin of 6% to 35% over the baselines. Notably, when using GPT-OSS-120B as a backbone, SR-Scientist achieves the highest overall performance, with an $Acc_{0.01}$ of 63.57% and an $Acc_{0.001}$ of 49.35%. In terms of performance across different subjects, with backbones such as Qwen3-Coder-480B, GLM-4.5-Air, and GPT-OSS-120B, SR-Scientist surpasses other LLM-based methods in all subjects. For Qwen3-Coder-30B, end-to-end RL training significantly improves its performance compared to the original model in all subjects, highlighting that the agent can evolve its abilities through its own experience. We provide further analysis of the LLM backbone for RL training in Appendix A.3.4 and the computation cost discussion in Appendix A.5.
@@
Figure 3: Noise robustness analysis. Qwen, GLM, and GPT represent Qwen3-Coder-480B, GLM-4.5-Air, and GPT-OSS-120B, respectively.
@@
Table 2: Symbolic accuracy (SA) of different methods. GLM, GPT, and Qwen represent GLM-4.5-Air, GPT-OSS-120B, and Qwen3-Coder-480B, respectively.
@@
Figure 2 shows the performance of the discovered equations on OOD data. For material science, the performance of all methods improves when shifting from ID to OOD data, while for other subjects, the methods exhibit varying trends. Among them, SR-Scientist still achieves the best performance on OOD data, demonstrating its strong generalization capabilities. The results of other models are presented in Appendix A.6 and show similar trends.
@@
Table 3: Ablation studies. In ‘w/o $T_{1}$’, the agent can not utilize the data analyzer tool. In ‘w/o experience’, the agent can not utilize the experience buffer and optimize from scratch for each iteration. GPT, Qwen, and GLM represent GPT-OSS-120B, Qwen3-Coder-480B, and GLM-4.5-Air, respectively.
Robustness to Noise
Symbolic Accuracy
4.3 Analysis
Ablation studies
Table 3 presents the results of our ablation studies on the data analysis and experience buffer components. The findings indicate that allowing the agent to perform data analysis is crucial for performance, particularly for the GPT-OSS-120B and Qwen3-Coder-480B models. Moreover, the experience buffer is essential for the continuous optimization of equations across iterations.
The Effect of the Turn Number

### v2

@@ Abstract
Recently, Large Language Models (LLMs) have been applied to scientific equation discovery, leveraging their embedded scientific knowledge for hypothesis generation. However, current methods typically confine LLMs to the role of an equation proposer within search algorithms like genetic programming. In this paper, we present SR-Scientist, a framework that elevates the LLM from a simple equation proposer to an autonomous AI scientist that writes code to analyze data, implements the equation as code, submits it for evaluation, and optimizes the equation based on experimental feedback. Specifically, we wrap the code interpreter into a set of tools for data analysis and equation evaluation. The agent is instructed to optimize the equation by utilizing these tools over a long horizon with minimal human-defined pipelines. Empirical results show that SR-Scientist outperforms baseline methods by an absolute margin of 6% to 35% on datasets covering four science disciplines. Additionally, we demonstrate our method’s robustness to noise, the generalization of the discovered equations to out-of-domain data, and their symbolic accuracy. Furthermore, we develop an end-to-end reinforcement learning framework to enhance the agent’s capabilities [^1] Code and data are available: https://github.com/GAIR-NLP/SR-Scientist.
@@ 1 Introduction
In the era of Agentic AI, Large Language Models (LLMs) have evolved from simple knowledge retrievers to agentic models capable of completing complex tasks by interacting with their environments (Schneider, 2025), as exemplified by products like Claude Code (Anthropic, 2025) and Gemini CLI (Google, 2025). These agents exhibit many characteristics of human scientists, such as engaging in long-horizon interaction with environmental feedback, operating with autonomy, and relying less on predefined pipelines (Newell et al., 1972). However, in most current work using LLMs for scientific discovery (Shojaee et al., 2025a; Novikov et al., 2025; Romera-Paredes et al., 2024), LLMs serve as static components within human-crafted pipelines, lacking the autonomy to generate and refine hypotheses through active environmental interaction. Building a scientific discovery framework around agentic models could therefore shift the paradigm from using LLMs as passive tools to empowering them as autonomous agents that drive the entire discovery lifecycle. In this paper, we focus on equation discovery, a fundamental task in science.
Mathematical equations are central to scientific progress, serving as concise and interpretable models of physical phenomena (Lemos et al., 2022; Batra et al., 2021; Hernandez et al., 2019). The task of data-driven equation discovery, also known as symbolic regression (SR), is an NP-hard problem due to its vast search space (Virgolin and Pissis, 2022). Traditional methods have relied on Genetic Programming (GP) for combinatorial search (Cranmer, 2023) or on deep neural networks trained on large-scale synthetic data for direct prediction (Kamienny et al., 2022; Biggio et al., 2021), which often suffer from efficiency and scalability issues. Recently, the research community has begun to embed LLMs into GP algorithms as equation proposers, leveraging their scientific prior knowledge to generate more effective hypotheses (Shojaee et al., 2025a; Grayeli et al., 2024). While this can enhance search efficiency, LLMs primarily serve as equation generators in a fixed pipeline, which lacks the autonomy to gain insight for equation design through tools.
@@
To validate the effectiveness of our framework, we evaluate SR-Scientist using 5 different models as backbone LLMs on datasets covering 4 science disciplines carefully designed to prevent the LLMs from relying on memorized equations from their training data (Shojaee et al., 2025b). Results demonstrate that SR-Scientist consistently outperforms state-of-the-art SR methods, with four of the models achieving an absolute margin of 6% to 35% over baselines. We also observe significant improvement after applying RL training. Additionally, we demonstrate our method’s robustness to noise, the generalization of the discovered equations to out-of-domain data, and their symbolic accuracy. Furthermore, our ablation studies highlight the importance of data analysis, the experience buffer, and long-horizon optimization.
@@ 2 Related Work
SR aims to identify an interpretable mathematical expression that accurately describes the relationship between variables in observed data (Makke and Chawla, 2024; Cava et al., 2021). Current popular methods can be categorized as follows: 1) GP-based Methods: This line of methods uses constrained representations, such as expression trees, to define a search space composed of symbolic operators, constants, and variables (Cranmer, 2023). Evolutionary or genetic programming is then applied to search for solutions. 2) Deep learning Methods: These methods leverage deep neural networks to learn a direct mapping from numerical observations to their underlying mathematical expressions through online (Petersen et al., 2021; Landajuela et al., 2022) or offline learning (Kamienny et al., 2022; Biggio et al., 2021). 3) LLM-based methods: Due to the efficiency and scalability issues of the above methods, the research community has integrated LLMs with traditional techniques like GP, leveraging the models’ extensive knowledge to generate more informed hypotheses (Shojaee et al., 2025a; Grayeli et al., 2024; Ma et al., 2024; Wang et al., 2025b). While promising, current hybrid methods still lack the direct analysis of observed data through tools to gain insights. Furthermore, the LLM’s behavior is often fixed, lacking the autonomy to decide its actions in a goal-oriented manner. Moreover, most work focuses on using LLMs for inference, without exploring how these models evolve their abilities through methods like RL.
@@ 3.2 Inference Framework
Tool Design
We provide a code interpreter as the primary tool, enabling the agent to interact with the data and validate hypotheses. Specifically, we wrap the code interpreter into two common tools for the tasks, data analyzer and equation evaluator, denoted as $T_{1}$ and $T_{2}$, respectively. For $T_{1}$, we link it to the observed data and include a code example in the prompt demonstrating how to access the data. Through this tool, the agent can write and execute diverse code to analyze the data, such as inspecting data samples, conducting statistical analysis, or analyzing the residuals between the predicted value and the true value [^3] Since we primarily use a text-only model that cannot accept figures as input, we explicitly instruct the model via prompts to avoid generating plots. However, the model backbone is orthogonal to our framework, which intrinsically supports such plot analysis patterns.. Notably, we do not constrain the code to these examples, allowing for diverse and emergent analysis patterns. For $T_{2}$, following the approach of Shojaee et al. (2025a), we design the tool to accept an equation skeleton with placeholders for constants in code format (See Figure 1 or 4 for the equation example). During the execution phase, the BFGS algorithm is implemented to optimize these constants and then report the performance of the equation. It also accepts equations with constants decided through other methods like data analysis. This tool prevents the agent from writing repetitive code for equation evaluation during the long-horizon exploration process.
@@ 3.3 Training Framework
Following the approach of Shojaee et al. (2025b), we employ a mixed strategy of rule-based and model-based data synthesis. For each scientific scenario covering multiple variables, we instruct an LLM to synthesize potential relationships between the variables, typically in the form of an equation skeleton. For each skeleton, we use the model to determine the values for the constants based on their physical significance and thereby construct the full equation. Once the complete equation is formulated, we define an appropriate range of values for its variables and synthesize the observed data accordingly. This dataset is then split into training and evaluation sets. The training data is accessible to the agent during rollouts, while the evaluation data is used to measure performance and calculate the reward.
@@
where $s_{\text{max}}$ represents the maximum MAPE for which a non-zero reward can be gained. Additionally, we explore other reward functions and present the discussion in Appendix C.2.
@@ 4.1 Setup
Due to the vast corpus that LLMs have been pretrained on, the evaluation dataset must be carefully curated to prevent LLMs from having memorized the equations. We evaluate our methods on the synthetic part of the LLM-SRBench (Shojaee et al., 2025b), denoted as LSR-Synth [^4] We also considered LSR-Transform (Shojaee et al., 2025b) as a candidate for the evaluation dataset. However, we found that it might still have memorization issues, as nearly 50% of the problems achieved a sufficiently small error after the first iteration.. It combines known terms in the underlying equation with synthetic, novel terms to create problems that go beyond memorization. Furthermore, all equations and visualizations of their generated data points were verified by two subject matter experts to ensure scientific rigor. It contains 129 problems, spanning four scientific domains: chemistry (36), biology (24), physics (44), and material science (25). For each problem, it has three types of datasets: a training set accessible to the SR method, an in-domain (ID) test set, and an out-of-domain (OOD) test set.
@@
We use accuracy-to-tolerance as our main metric. We find it serves as a more robust metric when aggregating the overall results covering multiple problems than others, like Normalized Mean Squared Error. It is also more challenging than other metrics like $R^{2}$. Please refer to Appendix D.1 for a detailed analysis. Given a desired tolerance threshold $\tau$, the metric calculates whether the predicted values and the ground truth values satisfy it as follows:
@@
We discard the 5% worst predictions to prevent outliers due to the presence of the max operator, following Biggio et al. (2021); Kamienny et al. (2022). Additionally, we present symbolic accuracy, which checks if the discovered equations are identical to the ground truth equations. We present the details for the symbolic accuracy calculation in Appendix D.6.
@@
We comprehensively compare our methods against several baseline methods. For methods without LLMs, we include GPLearn, E2E (Kamienny et al., 2022), NeSymReS (Biggio et al., 2021), DSR (Petersen et al., 2021), uDSR (Landajuela et al., 2022), and PySR (Cranmer, 2023). For the LLM-based methods, we include LLM-SR and LaSR for comparison. For methods without LLMs, we constrain the number of equation candidates for each problem to 100,000. For methods with LLMs, we constrain the LLM calls to 1,000, with at most 1,000 equation candidates from each problem. For SR-Scientist, we set the maximum number of turns per iteration to 25, with a total of 40 iterations. We present the configuration details of the baseline methods in Appendix A and the configuration of SR-Scientist in Appendix B. We evaluate our method with different models as backbone LLMs, including Qwen3-Coder-480B-A35B (Qwen, 2025), GLM-4.5-Air (Zeng et al., 2025a), GPT-OSS-120B (OpenAI, 2025), GPT-OSS-20B, and Qwen3-Coder-30B-A3B. Due to computation costs and training infrastructure limitations, we use only Qwen3-Coder-30B-A3B for RL training, while the other models are used for inference only. However, we also verify the effectiveness of RL training on other models and present the results in Appendix C.4. For each experiment, we repeat it three times and report the average value to reduce noise.
@@
For training data construction, we synthesize training data for the scientific scenarios in LSR-Synth. To prevent data contamination, two authors of this paper independently and carefully checked the equation skeletons to prevent duplicates and confirmed their agreement. Finally, we obtained 1024 problems for RL training. Please refer to Appendix C.1 for details. During rollout, we set the maximum number of turns to 20 and train for 60 steps. We set $s_{\text{max}}$ to 100% and $s_{\text{goal}}$ to 0.1% in the reward design. The detailed training configure can be found in the Appendix C.3.
@@ 4.2 Main Results
Table 1 presents the accuracy results. In terms of overall accuracy, SR-Scientist consistently outperforms the baseline methods, with four of the models achieving an absolute performance margin of 6% to 35% over the baselines. Notably, when using GPT-OSS-120B as a backbone, SR-Scientist achieves the highest overall performance, with an $Acc_{0.01}$ of 63.57% and an $Acc_{0.001}$ of 49.35%. In terms of performance across different subjects, with backbones such as Qwen3-Coder-480B, GLM-4.5-Air, and GPT-OSS-120B, SR-Scientist surpasses other LLM-based methods in all subjects. For Qwen3-Coder-30B, end-to-end RL training significantly improves its performance compared to the original model in all subjects, highlighting that the agent can evolve its abilities through its own experience. We provide further analysis of the LLM backbone for RL training in Appendix C.4 and the computation cost discussion in Appendix D.2.
@@
Figure 3: Noise robustness analysis on LSR-Synth. Qwen, GLM, and GPT represent Qwen3-Coder-480B, GLM-4.5-Air, and GPT-OSS-120B, respectively.
@@
Table 2: Symbolic accuracy (SA) of different methods on LSR-Synth. GLM, GPT, and Qwen represent GLM-4.5-Air, GPT-OSS-120B, and Qwen3-Coder-480B, respectively.
@@
Figure 2 shows the performance of the discovered equations on OOD data. For material science, the performance of all methods improves when shifting from ID to OOD data, while for other subjects, the methods exhibit varying trends. Among them, SR-Scientist still achieves the best performance on OOD data, demonstrating its strong generalization capabilities. The results of other models are presented in Appendix D.3 and show similar trends.
Robustness to Noise
Symbolic Accuracy
4.3 Analysis
Ablation studies
Table 3 presents the results of our ablation studies on the data analysis and experience buffer components. The findings indicate that allowing the agent to perform data analysis is crucial for performance, particularly for the GPT-OSS-120B and Qwen3-Coder-480B models. Moreover, the experience buffer plays an important role in the long-term memory of the agent, allowing it to utilize the experience of top-performing equations from previous iterations. Either directly removing the experience buffer or randomly sampling equations from it leads to a significant performance drop.
The Effect of the Turn Number
@@ 4.3 Analysis
w/o top-k 41.96 21.71
Table 3: Ablation studies. In ‘w/o $T_{1}$’, the agent can not utilize the data analyzer tool. In ‘w/o experience’, the agent can not utilize the experience buffer and optimize from scratch for each iteration. In ‘w/o top-k’, we randomly sample equations from the experience buffer instead of using top-k sampling. GPT, Qwen, and GLM represent GPT-OSS-120B, Qwen3-Coder-480B, and GLM-4.5-Air, respectively.
@@
5 Limitations and Future Directions
While SR-Scientist demonstrates significant advantages in equation discovery, we identify several limitations that open avenues for future research. First, in our main experiments, we only used text-only models. Expanding the model to incorporate multi-modal inputs and include plots in the analysis phase is a promising direction. Second, in noisy scenarios, although SR-Scientist performs best, it still suffers from a significant performance drop. Future work could explore incorporating common noise-handling strategies into the agent’s prompt to guide its data analysis. Third, since we only provide the top-performing equations to the agent in each iteration, the agent may still re-explore equations that previously performed poorly. Optimizing the agent’s memory system, such as by distilling high-level experience from the equations, is an interesting future direction.